OpenStudy (anonymous):

Which is a factor of 6x3y + 6xy − 12x2− 12? x + 2 xy − 1 x2 + 1 x2 − 2

OpenStudy (here_to_help15):

Do you need help or just the answer?

OpenStudy (here_to_help15):

You still there?

OpenStudy (here_to_help15):

6x3y + 6xy – 12x2– 12 =6xy(x^2+1)-12(x^2+1) =6(xy-2)(x^2+1)